# Information-Theoretic Bounds and Task-Centric Learning Complexity for Real-World Dynamic Nonlinear Systems

### v1

@@ Abstract
Index Terms:
Behavioral modeling, Complexity, Dynamic nonlinear systems, Inner product bounds, Learning complexity, Machine learning, Power amplifiers, Residual learning, Structured decomposition, System identification, Variance bounds.
@@ I Introduction
Our earlier work introduced an AI-based block-structured framework for modeling dynamic nonlinear systems, demonstrated on power amplifiers in communication systems [16]. A PA exemplifies a dynamic nonlinear system, with nonlinearities arising from static effects such as AM–AM and AM–PM conversion, and dynamic effects including thermal variation, bias drift, and memory [17]. The framework in [16] employs two blocks: a static nonlinear block and a neural network block. By isolating the dynamic residual and modeling only this component with a neural network, the method adapts to time-varying conditions while reducing training and inference complexity. This selective learning strategy effectively compensates for the nonlinear memory without modeling the entire distortion space. Experimental validation on a real PA showed superior performance over conventional approaches such as those in [18].
Unlike the classical progression where theoretical results precede experimental validation, this work originated from an empirical anomaly. Our apparent empirical success raised a deeper theoretical question:
@@ III Physical and Statistical Foundations for Modeling Real-World Dynamic Nonlinear Systems
This section develops results that formalize the physical and statistical constraints governing real-world dynamic nonlinear systems, spanning domains from engineering and control to finance, neuroscience, and astrophysics [1, 21, 22, 23]. While variance-based analysis is traditionally applied to power amplifier (PA)s, stock markets, and control systems, the underlying principles extend naturally to systems governed by nonlinear energy transport and bounded resource flow.
@@ V-A Variance as a Measure of Informational Uncertainty
This section develops a framework that connects the statistical variance of a target function with its mean-squared Lipschitz continuity and, ultimately, with the computational effort required to learn it. Variance is interpreted as a surrogate for informational uncertainty: a function with larger variance is more irregular, less predictable, and demands more information to approximate or compress within a prescribed error tolerance. While the formulation does not rely explicitly on Shannon entropy, it is consistent with the broader information-theoretic principle that unpredictability corresponds to informational richness. This perspective motivates the use of information-theoretic bounds in our framework for structured modeling and residual learning.
@@
Traditional complexity measures suchas VC dimension [26], Rademacher complexity [27], norm-based bounds [28], and kernel methods [29], quantify the expressive capacity of hypothesis classes but give limited attention to the structure of the target function itself. Our approach, in contrast, shifts focus to the intrinsic variability of the function to be learned. We establish that functions with higher variance, and therefore larger mean-squared Lipschitz constants, are inherently more difficult to learn under fixed error tolerance. This conclusion holds independently of the model class or learning algorithm, leading to a task-centric measure of difficulty grounded in the function’s structure.
@@
The conceptual foundation is reminiscent of Kolmogorov complexity [30], which characterizes an object by the length of its shortest description. Although Kolmogorov complexity is uncomputable in practice, it motivates our variance-based diagnostic, which is both tractable and measurable under finite data and noise. By quantifying a function’s irregularity through variance and mean-squared Lipschitz continuity, we obtain a practical and interpretable indicator of learnability.
@@
Our first theorem establishes that, under mild assumptions, the distortion component $d(x)$ of a structured nonlinear system possesses a strictly larger mean-squared Lipschitz constant than the residual $h(x)$, implying that the residual is smoother and less variable. The second theorem builds on this result and shows that smoother functions reduce learning complexity in four respects: parameter efficiency, convergence rate, generalization, and computational cost. These results provide theoretical justification for decomposing nonlinear systems into static and dynamic components, and for restricting learning to the residual dynamics.
@@
This insight also clarifies the success of our hybrid residual learning strategy in power amplifier linearization [16], where classical methods captured the static component $G$ and neural networks were applied only to the residual distortion $d$. The resulting improvements in performance and model compactness are consistent with the theoretical prediction that smoother residuals are easier to learn.
@@
To the best of our knowledge, this constitutes one of the first systematic efforts to link learning complexity to variance through the mean-squared Lipschitz constant. Classical Poincaré-type inequalities connect variance and Lipschitz continuity, but only for individual functions under specific measures. They do not extend naturally to function comparisons, such as between a distortion and its residual, since traditional Lipschitz constants measure worst-case fluctuations, whereas variance reflects average fluctuations. To address this mismatch, we introduce the mean-squared Lipschitz constant, defined as the expected squared gradient magnitude. This measure aligns more closely with variance and enables consistent comparison of complexity across functions.
@@
The resulting framework is architecture-independent, learning-agnostic, and broadly applicable in regression, control, and signal transformation. It complements existing tools such as PAC-Bayes bounds [31], Barron-type approximation results [32], and minimum description-length principles [33], by offering a measurable, task-level diagnostic of learning difficulty.
@@ V-C Implication of Complexity Reduction from Mean-Squared Lipschitz Dominance
The following theorem is formulated in a general fashion and is applicable to any setting where mean-squared Lipschitz dominance is observed between target functions. It provides a formal foundation for understanding how learning smoother functions—those with smaller mean-squared Lipschitz constants—leads to reduced learning complexity when using neural network approximators. This framework is independent of domain or application and thus can be employed broadly in signal processing, machine learning, and nonlinear system modeling wherever such dominance exists.
@@
Part 1: Parameter Efficiency.
@@
Part 2: Training Convergence.
@@
Part 4: Computational Cost.
@@
Hence, learning the smoother function $f_{2}$ incurs lower overall computational cost. ∎ This theorem rigorously establishes that learning a target function with a smaller mean-squared Lipschitz constant results in lower learning complexity, across multiple dimensions—parameter efficiency, convergence speed, generalization behavior, and computational cost. In the context of structured modeling of dynamic nonlinear systems, this result justifies a hybrid learning strategy: first approximate the static or relatively more predictable component $G$ using classical methods (e.g., polynomial fitting or nonlinear regression), and then use a neural network to model only the dynamic residual $d$, which is typically the more complex and relatively less predictable part.
@@ VI-A Measurement Setup
The experimental setup utilized a SKY66292-11 PA from Skyworks Inc., operating at 2.35 GHz with a 64-Quadrature Amplitude Modulation signal of 20 MHz bandwidth and 10.6 dB peak-to-average power ratio [41]. Refer [16], for additional details.
@@
While ARVTDNN required 25 input features, the proposed SRTDNN model used only 5, significantly reducing the input dimensionality. In both cases, a three-layer neural network was used comprising an input layer, one hidden layer, and an output layer. ARVTDNN achieved optimal performance with 25 hidden neurons, whereas SRTDNN used 50 neurons for improved residual learning. In SRTDNN, the static nonlinearity was first estimated via classical polynomial fitting, and the residual was then modeled by the neural network.
@@
These observations anchor the theoretical framework in experimental reality and validate the structured modeling paradigm as both computationally efficient and performance-effective. Moreover, the proposed variance-based analysis and memory indicators may offer useful priors for physics-informed machine learning (PIML) in scenarios where partial physical knowledge coexists with data-driven modeling. We also softly recommend applying classical mathematical tools to the relatively more predictable components, and reserving machine learning techniques for the relatively less predictable residuals. This strategy may enhance learnability and reduce computational burden—though, as our framework suggests, ideal learning remains near impossible due to inherent structural tradeoffs and behavioral uncertainty, mandating realistic performance thresholds.
@@ VII Conclusion
[…]
@@
Keysight Inc. has supported the research with donations of measurement equipment in [16]. Bulusu is deeply grateful for the profound knowledge imparted by Professor Emeritus of IIT Kanpur, Dr. Kalluri Ramalinga Sarma (Ph.D., Cornell University 1961), currently an adjunct professor at Mahindra University, through his insightful lectures in the areas of signals and systems, communication theory, and detection and estimation theory. The authors would like to thank Manoj Kumar Yadav (Ph.D., IIT Madras 2010), currently associate professor at Mahindra University, for reviewing the manuscript and identifying an important error in an earlier version of the proof of the Theorem IV.1.
@@ B-A Interpretation and Practical Relevance of the Weak Uncorrelatedness Lemma
Table III catalogs examples from engineered domains like PAs and optical fibers to biological, economic, and astrophysical systems, where static and dynamic components often evolve on distinct timescales with limited statistical coupling. Typically, the static part reflects memoryless, input-driven behavior (e.g., gain curves, equilibrium responses), while the dynamic part captures delayed or adaptive effects (e.g., memory, feedback, transients). The weak uncorrelatedness condition $|\mathbb{E}[(G-X)h]|\leq\epsilon$ thus defines a diagnostic uncertainty band—quantifying tolerable interaction between static and dynamic effects without losing analytical tractability.
@@
The Weak Uncorrelatedness Lemma player a formative role shaping our theoretical framework. This lemma initially served as an important assumption and revealed the plausibility of the critical condition $\theta+2\alpha>0$, which ultimately guided the finding of the DOB inequality. While now superseded in terms of theoretical necessity, the lemma retains interpretive value and reflects a recurring empirical pattern: across a variety of systems, static and dynamic components often exhibit weak statistical coupling. This observation motivates an open question—whether such decoupling reflects a deeper principle that prevents full correlation, such as that required to saturate the Cauchy–Schwarz bound. With the DOB Theorem establishing a lower bound, and empirical evidence from Table III suggesting a tight but nonzero interaction, we are led to the constrained interval:
@@
where $\epsilon_{0}$ represents a theoretical or empirical upper bound that remains to be identified. Clarifying the nature of this upper bound—whether system-specific or universal—remains an open challenge and an intriguing direction for future work.

### v2

@@ I Introduction
Our earlier work introduced an AI-based block-structured framework for modeling dynamic nonlinear systems, demonstrated on power amplifiers in communication systems [16]. A power amplifier (PA) exemplifies a dynamic nonlinear system, with nonlinearities arising from static effects such as AM–AM and AM–PM conversion, and dynamic effects including thermal variation, bias drift, and memory [17]. The framework in [16] employs two blocks: a static nonlinear block and a neural network block. By isolating the dynamic residual and modeling only this component with a neural network, the method adapts to time-varying conditions while reducing training and inference complexity. This selective learning strategy effectively compensates for the nonlinear memory without modeling the entire distortion space. Experimental validation on a real PA showed superior performance over conventional approaches such as those in [18].
Unlike the classical progression where theoretical results precede experimental validation, this work originated from an empirical observation in [16] that confirmed the abstract idea of separately learning static and dynamic effects. Our apparent empirical success raised a deeper theoretical question:
@@ III Physical and Statistical Foundations for Modeling Real-World Dynamic Nonlinear Systems
This section develops results that formalize the physical and statistical constraints governing real-world dynamic nonlinear systems, spanning domains from engineering and control to finance, neuroscience, and astrophysics [1, 21, 22, 23]. While variance-based analysis is traditionally applied to PAs, stock markets, and control systems, the underlying principles extend naturally to systems governed by nonlinear energy transport and bounded resource flow.
@@ V-A Variance as a Measure of Informational Uncertainty
In this section, we develop a framework connecting the variance of a target function with its mean-squared Lipschitz continuity and, ultimately, the computational effort required to learn it. Variance serves as a surrogate for informational uncertainty: functions with larger variance are more irregular and demand more information to approximate within a given error tolerance. Although our formulation does not invoke Shannon entropy, it is consistent with the principle that unpredictability corresponds to informational richness. This motivates the use of information-theoretic bounds in our framework for structured modeling and residual learning.
@@
Classical complexity measures such as VC dimension [26], Rademacher complexity [27], norm-based bounds [28], and kernel methods [29] emphasize hypothesis class capacity but largely ignore the structure of the target function. Our approach shifts focus to intrinsic variability: higher variance, and thus larger mean-squared Lipschitz constants, implies greater learning difficulty under fixed error tolerance. This task-centric view holds regardless of model class or algorithm.
@@
The idea recalls Kolmogorov complexity [30], which characterizes an object by the length of its shortest description. While uncomputable in practice, it motivates our variance-based diagnostic, which is tractable under finite data and noise. By quantifying irregularity through variance and mean-squared Lipschitz continuity, we obtain a measurable indicator of learnability.
@@
Our first theorem shows that, under mild assumptions, the distortion component $d(x)$ of a structured nonlinear system has a strictly larger mean-squared Lipschitz constant than the residual $h(x)$, implying smoother residuals. The second theorem demonstrates that smoother functions reduce learning complexity in terms of parameter efficiency, convergence, generalization, and computational cost. These results justify decomposing nonlinear systems into static and dynamic components, restricting learning to the residual.
@@
This also clarifies the success of our hybrid residual learning strategy in power amplifier linearization [16], where classical methods modeled the static part $G$ and neural networks handled the residual $d$. The observed gains in performance and compactness align with the prediction that smoother residuals are easier to learn.
@@
To our knowledge, this is among the first systematic efforts to link learning complexity to variance via mean-squared Lipschitz constants. Classical Poincaré-type inequalities connect variance and Lipschitz continuity for individual functions but do not extend to function comparisons. The mean-squared Lipschitz constant, defined as the expected squared gradient magnitude, resolves this by aligning more directly with variance and enabling consistent complexity comparisons.
@@
The framework is architecture-independent, learning-agnostic, and broadly applicable in regression, control, and signal transformation. It complements Probably approximately correct-Bayes bounds [31], Barron-type results [32], and minimum description-length principles [33], by providing a measurable, task-level diagnostic of learning difficulty.
@@ V-C Implication of Complexity Reduction from Mean-Squared Lipschitz Dominance
The following theorem is formulated in a general fashion and is applicable to any setting where mean-squared Lipschitz dominance is observed between target functions. It provides a formal foundation for understanding how learning smoother functions, those with smaller mean-squared Lipschitz constants, leads to reduced learning complexity when using neural network approximators. This framework is independent of domain or application and thus can be employed broadly in signal processing, machine learning, and nonlinear system modeling wherever such dominance exists.
@@
Part 1: Parameter Efficiency
@@
Part 2: Training Convergence
@@
Part 4: Computational Cost
@@
Hence, learning the smoother function $f_{2}$ incurs lower overall computational cost. ∎ This theorem rigorously establishes that learning a target function with a smaller mean-squared Lipschitz constant results in lower learning complexity, across multiple dimensions—parameter efficiency, convergence speed, generalization behavior, and computational cost. In the context of structured modeling of dynamic nonlinear systems, this result justifies a hybrid learning strategy: first approximate the static or relatively more predictable component $G$ using classical methods (e.g., polynomial fitting), and then use a neural network to model only the dynamic residual $d$, which is typically the more complex and relatively less predictable part.
@@ VI-A Measurement Setup
The experimental setup utilized a SKY66292-11 PA from Skyworks Inc., operating at 2.35 GHz with a 64-Quadrature Amplitude Modulation signal of 20 MHz bandwidth and 10.6 dB peak-to-average power ratio [41]. Refer [16], for additional details.The sampling rate of the baseband signal was 119.8 Msps, with a signal length of 1 ms. A simplified residual learning framework was employed: the static nonlinearity was modeled using polynomial fitting, and a shallow neural network was used to capture the dynamic residual.
@@
While ARVTDNN required 25 input features, the proposed SRTDNN model used only 5, significantly reducing the input dimensionality. In both cases, a three-layer neural network was used, comprising an input layer, one hidden layer, and an output layer. ARVTDNN achieved optimal performance with 25 hidden neurons, whereas SRTDNN used 50 neurons for improved residual learning. In SRTDNN, the static nonlinearity was first estimated via classical polynomial fitting, and the residual was then modeled by the neural network.
@@
These observations anchor the theoretical framework in experimental reality and validate the structured modeling paradigm as both computationally efficient and performance-effective. Moreover, the proposed variance-based analysis and memory indicators may offer useful priors for physics-informed machine learning (PIML) in scenarios where partial physical knowledge coexists with data-driven modeling. We also softly recommend applying classical mathematical tools to the relatively more predictable components, and reserving machine learning techniques for the relatively less predictable residuals. This strategy may enhance learnability and reduce computational burden, although, as our framework suggests, ideal learning remains near impossible due to inherent structural tradeoffs and behavioral uncertainty, mandating realistic performance thresholds.
@@ VII Conclusion
At its core are two behavioral indicators, $\theta$ and $\alpha$, which quantify static and dynamic effects. These lead to variance inequalities that reveal: (1) dynamic residuals are smoother than total distortion, and (2) memory in physical systems is inherently finite, a consequence derived from the First Law of Thermodynamics via the DOB Theorem, unlike classical thermodynamic bounds such as Landauer’s and Bekenstein’s, which rely on the Second Law.
@@
Keysight Inc. has supported the research with donations of measurement equipment in [16]. Bulusu is deeply grateful for the profound knowledge imparted by Professor Emeritus of IIT Kanpur, Dr. Kalluri Ramalinga Sarma (Ph.D., Cornell University 1961), currently an adjunct professor at Mahindra University, through his insightful lectures in the areas of signals and systems, communication theory, and detection and estimation theory. The authors would like to thank Dr. Manoj Kumar Yadav (Ph.D., IIT Madras 2010), currently an associate professor at Mahindra University, for reviewing the manuscript and identifying an important error in an earlier version of the proof of the Theorem IV.1.
@@ B-A Interpretation and Practical Relevance of the Weak Uncorrelatedness Lemma
Table III catalogs examples from engineered domains like PAs and optical fibers to biological, economic, and astrophysical systems, where static and dynamic components often evolve on distinct timescales with limited statistical coupling. Typically, the static part reflects memoryless, input-driven behavior (e.g., gain curves, equilibrium responses), while the dynamic part captures delayed or adaptive effects (e.g., memory, feedback, transients). The weak uncorrelatedness condition $|\mathbb{E}[(G-X)h]|\leq\epsilon$ thus defines a diagnostic uncertainty band, quantifying tolerable interaction between static and dynamic effects without losing analytical tractability.
@@
The Weak Uncorrelatedness Lemma played a formative role shaping our theoretical framework. This lemma initially served as an important assumption and revealed the plausibility of the critical condition $\theta+2\alpha>0$, which ultimately guided the finding of the DOB inequality. While now superseded in terms of theoretical necessity, the lemma retains interpretive value and reflects a recurring empirical pattern: across a variety of systems, static and dynamic components often exhibit weak statistical coupling. This observation motivates an open question, whether such decoupling reflects a deeper principle that prevents full correlation, such as that required to saturate the Cauchy–Schwarz bound. With the DOB Theorem establishing a lower bound, and empirical evidence from Table III suggesting a tight but nonzero interaction, we are led to the constrained interval:
@@
where $\epsilon_{0}$ represents a theoretical or empirical upper bound that remains to be identified. Clarifying the nature of this upper bound, whether system-specific or universal, remains an open challenge and an intriguing direction for future work.